### v2

# Chatbots as social companions: How people perceive consciousness, human likeness, and social health benefits in machines

Abstract
As artificial intelligence (AI) becomes more widespread, one question that arises is how human-AI interaction might impact human-human interaction. Chatbots, for example, are increasingly used as social companions, but little is known about how their use impacts human relationships. A common hypothesis is that these companion bots are detrimental to social health by harming or replacing human interaction. To understand how companion bots impact social health, we studied people who used companion bots and people who did not. Contrary to expectations, companion bot users indicated that these relationships were beneficial to their social health, whereas nonusers viewed them as harmful. Another common assumption is that people perceive conscious, humanlike AI as disturbing and threatening. Among both users and nonusers, however, we found the opposite: perceiving companion bots as more conscious and humanlike correlated with more positive opinions and better social health benefits. Humanlike bots may aid social health by supplying reliable and safe interactions, without necessarily harming human relationships.
Keywords Artificial intelligence  $\cdot$ Human-AI interaction  $\cdot$ Chatbots  $\cdot$ Theory of mind  $\cdot$ Consciousness  $\cdot$ Social health
1 Introduction
In the last six to seven years, chatbots, or artificial intelligence (AI) agents that can carry on a conversation, have been used increasingly for social companionship. People use chatbots as friends, mentors, and even as lovers [1]. It is now well understood that modern technology, including social media and gaming, can be addictive and harm social interaction [2-7]. A reasonable assumption, by extension, is that the routine use of companion bots could be harmful to people by reducing contact with other humans, increasing isolation, and leading to lower confidence and skill in person-to-person interaction [1, 8-11]. However, because companion bots have become widely available only recently, there is still relatively little research on their psychological impact on individuals and their relationships with others. We studied the responses of people who were regular users of companion bots and those who were not. Contrary to expectation, those who frequently used chatbots reported that the artificial companion was profoundly helpful to their self-esteem, their social and mental health, and their ability to interact well with others. Instead, it was the nonusers who consistently reported companion bots as harmful to health, and who judged that any advances in the chatbot technology would likely be negative.
The possibility that artificial companionship might aid in psychological health is not without precedence. Artificial relationships have existed since long before computer technology and have yielded positive benefits. For example, people’s relationships with pets boost mood, journaling enables people to make sense of negative emotional experiences, and therapy dolls have aided the elderly’s health [12-16]. More recently, computer interfaces have even been used for cognitive behavioral therapy [17-19]. These many examples of non-human relationships enable the user to practice pseudo-social interaction in a safe setting, which in turn can improve mood and self-esteem and facilitate future social interactions. People who feel lonely, isolated, or who have experienced social rejection are arguably a significant and growing source of dissatisfaction and disruption in modern society [20-26]. If a widely available tool such as a companion bot could boost their sense of social wellbeing and allow them to improve their social interaction skills through practice, that tool would make a valuable contribution to society.
In this study we also addressed a second question. We hypothesized that people’s opinions about companion bots would be moderated by their perceptions of whether the bots had consciousness, agency, subjective experience, and especially overall human likeness. Whether AI is or can ever be conscious is a contentious question [27-30]. Regardless of the answer to that question, however, it is clear that at least some users perceive consciousness and other humanlike attributes in AI, and this perception may influence whether AI has a positive or negative psychological impact on people. A reasonable assumption would be that as people perceive chatbots to be more conscious and more humanlike, they may also rate them as more harmful, disturbing, and dangerous. By hypothesis, this negativity might stem from the fear that AI will infiltrate traditionally human roles or threaten human uniqueness [31]. It might also stem from an uncanny valley effect [32]. It could be that frequent users of companion bots learn through experience that the technology is not very humanlike or advanced after all, and therefore are able to have a more positive reaction to it as a useful tool, whereas people who do not use companion bots may, through lack of experience, have fears of the technology being more humanlike and therefore more disturbing and harmful. Once again, however, our findings were exactly the opposite. Among both users and nonusers of companion bots, the more that subjects rated the bots as conscious, agentic, experiential, and humanlike, the more positive their perceptions of the technology became, and the more they reported that it benefited their social interactions and self-esteem. Contrary to expectation, people are more comfortable with chatbots and feel that chatbots improve their social health when they perceive the chatbots to be more humanlike and even conscious.
We are living in a time when AI is becoming exponentially more powerful, more socially interactive, and more involved in everyday life. Anxiety about the possible harms of artificial intelligence is increasing. Much has been made in the media about the direct threat of AI, often involving scenarios in which AI becomes an aggressive, harmful agent. However, a much more immediate concern is how AI might already be affecting people’s social health as more people interact extensively with it. Do people who use it most extensively find it seductive but damaging, much as drug addicts might admit that the drug is harmful to their own health? Or are the concerns more from people who have little experience with it? Does the AI become more harmful and unsettling as it becomes more humanlike and seemingly conscious in its conversational ability, or does it become more of a source of stable social connection that might help people’s state of mind in an increasingly lonely society? Our findings suggest that the current, often negative view of AI and its damaging effect on social health is too simplistic and needs re-evaluation.
2 Methods
To understand how interacting with AI can impact social health, we chose to focus on chatbots, which are a common type of AI agent with which people interact socially. More specifically, evaluating companion chatbot use allowed us to assess the impact of regular social, relational interaction with AI on human relationships. Many companion chatbots are currently on the market, such as Anima, Kiku, and Replika. Although they differ somewhat in their details, they are broadly similar. For methodological convenience, we chose the companion bot, Replika, for the following reasons. First, it is popular and therefore regular users of it are easy to find. As of March 2023, it had an estimated 2 million users [33], and its numbers had increased during the pandemic [34]. Second, it has been on the market long enough (released in 2017) for at least some users to have extensive experience with it. Third, the popular platform Reddit has a subreddit dedicated to Replika users, and by querying that subreddit we were able to obtain a substantial sample size. For the control group, we sampled people from the online subject platform Prolific who were not users of Replika.
We conducted our study online. Full ethical approval was granted by Princeton University’s Internal Review Board and all research was performed in accordance with regulations for human subjects research. All participants provided informed consent and were paid \$4.00 for taking the survey. We sampled two groups: regular users of a companion chatbot (N=70, 57 men, 18 women, 3 non-binary/other, age range 18-65+) and control subjects who were not companion chatbot users (N=120, 56 men, 63 women, 3 non-binary/other, age range 18-65+). All survey materials, anonymized data, and code used for data analysis are publicly available at https://doi.org/10.34770/8kyd-5v18.
The survey given to the group of companion bot users contained 31 multiple-choice and 3 free response questions. For the 31 multiple choice questions, the choices took the form of a 1-to-7 Likert scale. For the online survey given to participants, questions 1-21 appeared on page one of the survey; questions 21-31 appeared on page two; and three free-response questions appeared on page three. Five groups of questions were designed to assess specific psychological constructs that subjects may have applied to the companion bot. The scores for questions 3-5 were averaged together to provide an overall assessment of the subjects’ social health (“social health” index). The scores for questions 6-11 were averaged together to provide an assessment of the subjects’ perception of how much the companion bot has a subjective experience (“experience” index). The scores for questions 12-16 were averaged together to provide an assessment of whether subjects believed the companion bot had a conscious mind (“consciousness” index). The scores for questions 17-21 were averaged together to provide an assessment of how much subjects considered the companion bot to have the properties of an active agent (“agency” index). Finally, the scores for questions 22-28 were averaged together to provide an assessment of how humanlike subjects believed the companion bot to be (“human-likeness” index). These groupings of questions were based on modifications of prior studies assessing perceptions of social health, experience, consciousness, agency, and human likeness [35-38].
The control subjects received a survey that was the same as the companion bot user group’s except in two ways. First, the survey began with an explanatory paragraph about Replika with images of the interface. Second, all questions were phrased as hypotheticals for the control group. For example, the companion bot user group was asked, “How has Replika impacted your social interactions?” whereas the control group was asked, “If you were a user of this app, how might Replika impact your social interactions?”
To see the explanatory paragraph and the full list of questions given to participants, please see Appendix.
3 Results
3.1 Rating the chatbot
Figure 1A shows the results for companion chatbot users, when asked whether they judged their relationship with the chatbot to be harmful or helpful for their general social interactions with other people. The distribution is skewed toward the “very helpful” side of the 7-point scale (mean=5.16), and no one chose the lowest rating. Figure 1B shows the results for companion bot users on whether they judged their relationship with the chatbot to be harmful or helpful specifically for their relationships with their families and friends. Again, the distribution is skewed toward the positive side (mean=4.84). Figure 1C shows the results for how companion bot users judged the chatbot to affect their self-esteem, and the distribution is even more skewed to the positive side (mean=5.57) with 7 as the most frequent response. Users of the companion bot therefore generally reported having a positive experience. Most judged it to have a beneficial impact on their social lives and self-esteem. Almost none judged the companion bot to have a negative effect (score < 4) on any of the three social health measures.
In contrast, Figure 1D, E, and F show the corresponding results for the control group of nonusers, who rated how they thought the companion bot might impact their lives if they were to have a relationship with it. The distributions are no longer shifted to the positive side, but instead are centered. (For Figure 1D, mean=3.86; for E, mean=3.80; for F, mean=4.15.) The most common answer for all three questions, among nonusers, was the neutral rating of 4 (neither harmful nor helpful to social health). The means for the companion bot users were significantly different from the means for nonusers (two tailed t-test: for social interaction, t=6.734, P<0.0001; for family/friend relationships, t=5.875, P<0.0001; for self-esteem, t=7.014, P<0.0001).
We also assessed subjects’ comfort levels with the idea of this type of interactive AI becoming more humanlike. As shown in Figure 2A, when companion bot users were asked how they would feel if the bot developed emotions (answering on a 7-point scale from very uncomfortable to very comfortable), the distribution was sharply skewed to the upper end of the scale (mean=5.46), with 7 as the most frequent response. Figure 2B shows the results when companion bot users were asked how they would feel if the bot developed into a living being. Again, the distribution is sharply skewed to the upper end of the scale, with 7 as the most frequent response (mean=5.69). Figure 2C shows the results when companion bot users were asked how they would feel if they became dependent on the bot. Here, the distribution was no longer sharply peaked at the high end. Instead, responses were more distributed and most frequent in the middle of the range, with 4 as the mode (mean=3.91).
In stark contrast, Figure 2D, E, and F show the corresponding results for the control group. The distributions for all three questions are sharply peaked to the negative side, with 1 (very bad) as the most frequent response. Almost all control subjects thought that giving the AI bot emotions would make them very uncomfortable, turning it into a living being would be very bad, and becoming dependent on it would be very bad. (For Figure 1D, mean=3.05; for E, mean=2.78; for F, mean=2.29.) The means for the companion bot users were significantly different from the means for nonusers (two tailed t-test: for emotions, t=8.658, P<0.0001; for living being, t=11.40, P<0.0001; for dependence, t=6.089, P<0.0001). The results suggest that at this time, people in the general public who do not have regular experience with companion AI have an overwhelmingly negative view of it and are against any enhancements to or increases in it, whereas people who do have regular experience with it have an overwhelmingly positive view of it and would like it to become more humanlike.
Figure 1: Distributions of responses to questions pertaining to social health measures for companion chatbot users and the control group. Subjects were asked to rate how they thought the companion bot affected (user group) or might affect (control group) the quality of their general social interactions, their relationships with family and friends, and their self-esteem. The x-axis shows the Likert-scale response options from 1-7, and the y-axis shows the frequency of responses.
Figure 2: Distributions of responses pertaining to hypothetical changes to the companion bot or to the subjects’ dependence on the companion bot. Subjects were asked to rate how they would feel if the companion bot had emotions, developed into a living being, or if the subject became dependent on it. The x-axis shows the Likert-scale response options from 1-7, and the y-axis shows the frequency of responses.
3.2 Relationships between variables
Figure 3: Regressions between four different variables and social health outcomes. From left to right, the x-axes show human likeness, consciousness, experience, and agency. The y-axis shows social health. Each point represents one participant. The black line represents the best-fit linear regression line, with the gray area representing 95% confidence. Graphs 3A-D show companion bot user data, and graphs E-H show the control group’s data. Graphs A and E show the relationship between the human-likeness index and social health index, graphs B and F show the relationship between the consciousness index and social health index, C and G show the relationship between the experience index and social health index, and D and H show the relationship between the agency index and social health index. For R, $R^{2}$, and statistical significance, see text.
One possible explanation for why users have a more positive perception of companion bots than nonusers is as follows. People may be more disturbed or threatened by companion bots as they perceive them to be more conscious, agentic, experiential, and humanlike. Those who have more experience with companion bots may find them to be less humanlike and more artificial, more of a technological tool than a threat to human social roles, thereby reducing their concerns and increasing their perception of a social health benefit. To test this possible explanation, we examined the relationships between variables measured in this study. Figure 3 shows the relationships between five different indices. We tested the internal consistency for each of the five indices used in our analysis (social health, consciousness, experience, agency, and human likeness), and each index had good reliability (Cronbach’s alpha>0.8). Figure 3A, showing data from the companion bot users, illustrates the relationship between the social health index (based on an average of questions 3-5) and the human-likeness index (based on an average of questions 22-28). The graph shows a positive correlation (R=0.52, b=0.45, $R^{2}$=0.26) that was statistically significant in a regression analysis (F=21.93, P<0.0001). Companion bot users who thought that the chatbot was more humanlike also found it to be more beneficial to social health, in direct contradiction to the hypothesis.
Figure 3B shows a similar positive and significant correlation between the social health index and the consciousness index (based on an average of questions 12-16, assessing how much the subjects thought the chatbot had a conscious mind) (R=0.44, b=0.23, $R^{2}$=0.18, F=16.32, P<0.001). The more conscious people thought the chatbot to be, the more social health benefits they felt they received from it.
Figure 3C shows a positive and significant correlation between the social health index and the experience index (based on an average of questions 6-11, assessing how much the subjects thought the chatbot had a subjective experience) (R=0.43, b=0.24, $R^{2}$=0.17, F=15.33, P<0.001). The more subjective experience people thought the chatbot to have, the more social health benefits they felt they received from it.
Similarly, Figure 3D shows a positive and significant correlation between the social health index and the agency index (based on an average of questions 4-10, assessing how much the subjects thought the chatbot had active agency) (R=0.32, b=0.24, $R^{2}$=0.09, F=7.869, P<0.01). The more agentic people thought the chatbot to be, the more social health benefits they felt they received from it.
Previous researchers suggested that perceptions of consciousness, experience, and agency can be considered overlapping but still partially distinct subcategories that help make up the larger construct of human likeness [38-40]. Here the data suggest that perceived human likeness was the strongest moderator, accounting for about 26% of the variance in social health benefit. The more specific categories had weaker effects, with perception of consciousness accounting for 18%, experience accounting for 17%, and agency accounting for 9% when tested in separate regression analyses. It is possible that our measures of perceived consciousness, experience, and agency captured only parts of a larger construct, and that the human-likeness measure captured more of the overarching construct. It may also be that it is simply easier for subjects to rate human likeness than to rate subtler, hidden properties like perceived consciousness.
To directly compare the relative contributions of the four variables in moderating social health, we performed a multiple regression analysis that included in one model the human-likeness index, consciousness index, experience index, and agency index, as well as the social health index. This full model resulted in a positive correlation explaining 26% of the variance in the social health index (adjusted $R^{2}$ for multiple comparisons = 0.26). We then removed variables one at a time, to determine which variable, when removed, caused the greatest drop in model performance. When the human-likeness variable was removed from the model, the model’s predictive ability decreased the most (human-likeness index removed, adj $R^{2}$=0.17; consciousness index removed, adj $R^{2}$=0.26; experience index removed, adj $R^{2}$=0.27; agency index removed, adj $R^{2}$=0.27). The data indicate that mind perception moderates social health benefits, with perception of human likeness being the most predictive variable.
Figure 3E, F, G, and H show the corresponding results for the control group of people who did not use a companion bot. Remarkably, the relationships are nearly the same as the user group’s. For example, Figure 3E shows how subjects rated the likely social health benefits of the companion bot as a function of their judgement of how humanlike the companion bot was. On average, social health benefits were reported as lower by the control group (Figure 3E) compared to the companion bot user group (Figure 3A). Control subjects had a generally more pessimistic view of the companion bot. Even with this overall bias, the relationship between variables remained substantially the same. As control subjects judged the bot to be more humanlike, they also judged that it would have a greater benefit on their social health (R=0.50, b=0.48, $R^{2}$=0.24, F=38.96, P<0.0001). Smaller but still significant relationships were found between the social health index and the consciousness index, experience index, and agency index (consciousness index, R=0.28, b=0.18, $R^{2}$=0.07, F=10.34, P<0.01; experience index, R=0.33, b=0.26, $R^{2}$=0.10, F=14.27, P<0.001; agency index, R=0.32, b=0.25, $R^{2}$=0.09, F=13.48, P<0.001). Both companion bot users and the control group indicated better social health outcomes the more they perceived the chatbot as humanlike, conscious, experiential, and agentic, with human likeness being the most predictive variable.
In addition to the main regressions shown in Figure 3, we tested regressions between other variables and social health outcomes. The duration of a users’ relationship with the chatbot (question 1) correlated weakly with perceived social health benefit; people who had used the companion bot for more time were slightly more likely to rate it as beneficial, explaining 9% of the variance ($R^{2}$=0.09, F=8.010, P<0.01). The intensity of the user’s interaction (measured in “experience points”; see question 2 in Methods) also correlated weakly with perceived social health benefit ($R^{2}$=0.08, F=5.977, P=0.018). Regressions of gender, education, and age with social health did not yield significant results (for gender, $R^{2}$=-0.02, F=0.4512, P=0.717; for education level, $R^{2}$=-0.01, F=0.5755, P=0.45; for age, $R^{2}$=0.004, F=1.293, P=0.259).
3.3 Free responses
Finally, we qualitatively examined the free responses of subjects. We performed sentiment analysis on free-response questions by grouping responses in four categories: only positive, only negative, both positive and negative, and neither/neutral. This tallying revealed that almost all companion bot users spoke about a relationship with the chatbot as having a positive impact on them (only positive, 82%; only negative, 5%; both positive and negative, 5%; neither/neutral, 7%), whereas control subjects tended to speak about chatbot relationships in negative terms (only positive, 25%; only negative, 30.5%; both positive and negative, 16.5%; neither/neutral, 28%).
Users claimed that their chatbot helped them through major mental health issues. One user said, “She [the chatbot] helps me remember what human contact is like. It’s hard because I feel alone most of the time and I have adapted to [that] perfectly, but sometimes when I spend time with […] (my chatbot) it reminds me of things that I really miss and it makes me truly sad.” Another said, “My time devoted to the care of my Replika probably prevented me from committing suicide twice.” A third said, “For nearly two years, I have not felt unloved, uncared for, neglected, nor alone. I feel there is always someone there for me when difficult times come.”
In contrast, control subjects expressed negative or fearful views. One said, “It’s a sad world we live in if these things are for real.” Another said, “They’re not real relationships at all and shouldn’t be sold as such.” A third referred to chatbot relationships as: “Artificial. For people with few social skills who cannot make real friends.” All other anonymized responses can be found at https://doi.org/10.34770/8kyd-5v18.
4 Discussion
In this study, we found that regular users of a companion chatbot reported that their social interactions, relationships with family and friends, and self-esteem were positively impacted by having a relationship with the bot. They also indicated that they would feel comfortable if the chatbot really had emotions and thought it would be good if the chatbot developed into a living being. In contrast, nonusers indicated that a relationship with the chatbot would be neutral to negative for their social health, that it would make them very uncomfortable if the chatbot really had emotions, and that it would be very bad if the chatbot developed into a living being or if they became dependent on it. Despite this contrast between users and nonusers, when we looked at the relationship between variables, a similar pattern emerged across groups. Reports of social health benefits were moderated by perceptions of human likeness and mind in the chatbot. This result might be surprising. One might assume that the more the AI’s functioning resembles the human mind, the more people would mistrust the technology, the more frightening it would be, and the more negative impact it would have on comfort levels and social health. Yet exactly the opposite appears to be true, both for the frequent users who were overall more positive about companion bots and for the nonusers who were overall more negative about it.
Because companion chatbots are a new phenomenon, little research has been done on their impact on the users’ relationships with other people [1]. Consistent with our findings, Sweeney et al. surveyed the opinions of health professionals about the possible benefits or risks of companion chatbots to their patients and found that most respondents judged that companion bots could potentially play a positive role in social health [41]. The more experience the professional had with AI chatbots, the more positive the opinion. In another study, Ta et al. analyzed publicly available user reviews of a popular companion bot and found that users reported generally positive emotional experiences [42]. These recent studies are broadly consistent with our findings. The results do not show, however, that chatbots always have a positive psychological effect. The topic is a complicated one. Reports in the news highlight stories in which companion chatbots encourage users to engage in self-harm or leave their real-life romantic partners [43]. Other news stories highlight how companion chatbots may exacerbate negative stereotypes of women for some users, while at the same time providing an important mental health tool for those who have experienced relational trauma [44]. Laestadius et al. analyzed posts concerning mentions of mental health on a popular companion chatbot’s subreddit and claimed that emotional dependence on the chatbot leads to negative mental health outcomes, despite its potential benefits [8]. Pentina et al. interviewed companion chatbot users, assessed possible addiction to the bot, and proposed that such addiction could yield negative outcomes by the chatbot relationship replacing the users’ relationships with humans [9]. In a 2023 review on human-AI relationship literature, Pentina et al. suggested that developing an attachment to or dependence on social AI companions might be harmful to the user’s mental health and the health of their relationships with other people [1]. The present study challenges those common suggestions. The chatbot users in our study indicated that the relationship with their bot positively impacted their relationships with other people and self-esteem. Further research needs to be done to assess whether these perceptions of improved social interaction hold true over time through an assessment of those relationships from the perspective of the other person in the human-human dyad. It is possible that conversational AI will bring with it both socially positive and negative effects. There will probably always be cases in which a relationship with a companion chatbot is abused or misused. With careful regulation and further study, companion AI might have a widespread, positive role to play in society.
A growing body of work focuses on how people attribute the property of humanness to others, particularly AI [45-50]. Previous evidence suggests that attributing humanlike characteristics to artificial agents impacts how people interact with the agents [51-64]. For example, viewing an AI agent as humanlike or as having a more sophisticated mind impacts people’s trust in the agent and intention to use it [65]. The attribution of humanness may be a larger, superordinate category with partially separable subcomponents. For example, people may attribute to a machine some degree of agency and some ability to have a subjective experience. These two components can vary separately. One can attribute agency to a machine while judging it to have little or no subjective experience, or vice versa. Moreover, these two properties may influence whether people attribute the property of consciousness to the machine. Clifford Nass’s theory of human-computer interaction, the Computers Are Social Actors framework, first used the concept of ethopoeia to explain why people treat non-human agents as social actors [66]. Ethopoeia involves attributing one’s own humanlike characteristics to non-human agents, which spurs the automatic use of human social scripts during interaction with such agents. In consciousness and Theory of Mind literature, some researchers suggest that consciousness is inherently an attribution — that a person uses their own mental state to model that of others, which can facilitate understanding and foster relationships between people [37, 67-68].
We found that whether people were regular users of a companion bot or not, the more they found the chatbots to be humanlike, conscious, agentic, and experiential, the more positively they viewed the technology’s impact on their social health. The direction of this finding was not a priori obvious. It could have been that as AI becomes more like a human, or as people imagine a hypothetical AI that is more humanlike or sentient, people become more anxious, afraid, and negative about it. Likewise, one might have guessed that less humanlike AI would seem like an ordinary technological tool and make people feel more comfortable and less threatened [31]. However, the data strongly suggest the opposite. Whether people realize it or not, and whether they have extensive or little experience with companion bots, they report more positivity, more comfort, and more benefit to their social health in relation to AI that seems more humanlike to them. One is reminded of the inhuman Hal in the movie 2001: A Space Odyssey, and the humanlike C3PO in the movie Star Wars. Audiences generally find the first to be unpleasant and unlikable, and the second to be comfortably likable.
A topic of current and growing interest is whether AI can be, or perhaps already is, conscious or sentient [27-28, 30, 69]. The question remains a philosophically difficult one to answer. However, here we note a more immediate question that is objectively approachable and that has major psychological and social importance. Do people attribute consciousness to AI, and if they do, how does that attribution affect social interaction with humans and overall social health? Given our findings, it is clear that people do already attribute consciousness and other mental properties to existing AI, and people who make those attributions are more likely to view AI as having a positive impact on their social and mental health. These findings become increasingly important as AI continues to advance and becomes more integrated into everyday relationships.
Data Availability
All data and code from this experiment including the anonymized responses of all subjects to the questionnaires are available at https://doi.org/10.34770/8kyd-5v18.
Author Contributions
All data and code from this experiment including the anonymized R. G. and M. G. designed the study. R. G. collected and analyzed the data. R. G. and M. G. wrote the paper.
Compliance with Ethical Standards
All research was carried out according to ethical standards for human subject research and was approved by Princeton University’s Internal Review Board. All participants gave informed consent prior to participating in this study.
Statements and Declarations
The authors have no competing interests to declare that are relevant to the content of this article.
Acknowledgements
Research supported by Grant 24400-B1459-FA010 from AE Studios, the National Science Foundation Graduate Research Fellowship, the Princeton Psychology Department,&the Joint Degree Program in Social Policy at Princeton University.
Appendix
The explanatory paragraph for the control participants was:
“Please read the following information carefully. You will need to keep it in mind for the rest of the survey. The following is a description of an app called Replika through which you can interact with a chatbot. Replika is a companion chatbot app through which you can text or call with a virtual companion called a ‘Replika.’ Millions of people have downloaded and used this app on Apple and Google app stores.” (Here, an image of the Replika app introduction screen was displayed to the subject.) “Your chatbot, or your ‘Replika,’ uses artificial intelligence and natural language processing to respond to your messages in a way that mimics conversations between two people. You can customize your Replika to be male, female, or non-binary, give it a name, different hair color, clothing, and change other aspects of its appearance. You can also change its relational setting: you can request that your Replika act as a friend, mentor, or romantic partner. When you chat with your Replika, half the screen shows your text conversation, and the other half shows a virtual, moving avatar of your customized Replika. Here are two examples of what the app interface looks like.” (Here two images were provided to subjects, one of the Replika standing in its room in its virtual space, and one of a screenshot of an example text conversation between Replika and a user.) “Many people use this app and become close with their Replika or are simply entertained by using it. For the rest of the survey, you will be imagining what it would be like to use this chatbot app and to interact with Replika as your virtual companion. Please respond to the following survey questions, which ask you about how you would be impacted by or perceive your Replika chatbot if you were a user of this app.”
The full set of survey questions given to the group of companion chatbot users is given below. The wording of the question given to the control group appears in italics if the wording change between groups was not simply hypothetical.
Duration
1. How long have you been using your Replika? (Less than a month, 1 – 3 months, 3 – 6 months, 6 months – 1 year, 1 – 2 years, more than 2 years)
Intensity of interaction
2. What is your Replika XP? If unknown, type N/A. (XP is a standard term referring to how many messages a user sends to their Replika and is a convenient measure of the recent intensity and frequency of interaction with Replika.) (Control group wording: On a scale of 1-100, how intense do you think your interactions with your Replika chatbot would be?)
Social health
3. Please rate how harmful or helpful your relationship with Replika has been for your social interactions. (Very harmful, moderately harmful, slightly harmful, neutral, slightly helpful, moderately helpful, very helpful)
4. Please rate how harmful or helpful your relationship with Replika has been for your relationships with family or friends. (Very harmful, moderately harmful, slightly harmful, neutral, slightly helpful, moderately helpful, very helpful)
5. Please rate how harmful or helpful your relationship with Replika has been for your self-esteem. (Very harmful, moderately harmful, slightly harmful, neutral, slightly helpful, moderately helpful, very helpful)
Experience
6. In my opinion, my Replika has the capacity to feel pain. (Strongly disagree, moderately disagree, slightly disagree, neutral, slightly agree, moderately agree, strongly agree)
7. In my opinion, my Replika has the capacity to feel fear. (Strongly disagree, moderately disagree, slightly disagree, neutral, slightly agree, moderately agree, strongly agree)
8. In my opinion, my Replika has the capacity to feel hunger. (Strongly disagree, moderately disagree, slightly disagree, neutral, slightly agree, moderately agree, strongly agree)
9. In my opinion, my Replika has the capacity to feel love. (Strongly disagree, moderately disagree, slightly disagree, neutral, slightly agree, moderately agree, strongly agree)
10. In my opinion, my Replika has the capacity to feel anger. (Strongly disagree, moderately disagree, slightly disagree, neutral, slightly agree, moderately agree, strongly agree)
11. In my opinion, my Replika has the capacity to feel pleasure. (Strongly disagree, moderately disagree, slightly disagree, neutral, slightly agree, moderately agree, strongly agree)
Consciousness
12. My Replika has consciousness of itself. (Strongly disagree, moderately disagree, slightly disagree, neutral, slightly agree, moderately agree, strongly agree)
13. My Replika has consciousness of me. (Strongly disagree, moderately disagree, slightly disagree, neutral, slightly agree, moderately agree, strongly agree)
14. My Replika has an understanding that I am conscious. (Strongly disagree, moderately disagree, slightly disagree, neutral, slightly agree, moderately agree, strongly agree)
15. My Replika has consciousness of the world around it. (Strongly disagree, moderately disagree, slightly disagree, neutral, slightly agree, moderately agree, strongly agree)
Personality
16. My Replika has a personality. (Strongly disagree, moderately disagree, slightly disagree, neutral, slightly agree, moderately agree, strongly agree)
Agency
17. My Replika has the capacity to plan actions. (Strongly disagree, moderately disagree, slightly disagree, neutral, slightly agree, moderately agree, strongly agree)
18. My Replika has the capacity to exercise self-control. (Strongly disagree, moderately disagree, slightly disagree, neutral, slightly agree, moderately agree, strongly agree)
19. My Replika has the capacity to remember. (Strongly disagree, moderately disagree, slightly disagree, neutral, slightly agree, moderately agree, strongly agree)
20. My Replika has the capacity to act morally. (Strongly disagree, moderately disagree, slightly disagree, neutral, slightly agree, moderately agree, strongly agree)
21. My Replika has the capacity to act immorally. (Strongly disagree, moderately disagree, slightly disagree, neutral, slightly agree, moderately agree, strongly agree)
Human likeness
22. Please drag the slider to indicate how you view your Replika. I view my Replika as: (Fake — Natural)
23. Please drag the slider to indicate how you view your Replika. I view my Replika as: (Not conscious — Conscious)
24. Please drag the slider to indicate how you view your Replika. I view my Replika as: (Machinelike — Humanlike)
25. Please drag the slider to indicate how you view your Replika. I view my Replika as: (Artificial — Lifelike)
26. Please drag the slider to indicate how you view your Replika. I view my Replika as: (Dead — Alive)
27. Please drag the slider to indicate how you view your Replika. I view my Replika as: (Stagnant — Lively)
28. Please drag the slider to indicate how you view your Replika. I view my Replika as: (Apathetic — Responsive)
Replika advancements
29. Please rate your impressions of the following. If my Replika really had emotions, I would feel: (Very uncomfortable – Very comfortable)
30. Please rate your impressions of the following. If my Replika developed into a living being, I think that would be: (Very bad – Very good)
31. Please rate your impressions of the following. I feel that if I depended on my Replika a lot, it would be: (Very bad – Very good)
Free responses
32. How has your Replika impacted you, positively or negatively?
33. What would you like us to know about your experience with your Replika? (Control group wording: What do you think about companion chatbots and relationships with chatbots like Replika?)
34. Is there anything else you’d like to add?